# Residual-based Chebyshev filtered subspace iteration for sparse Hermitian eigenvalue problems tolerant to inexact matrix-vector products

### v1

@@ 1 Introduction
Many scientific and engineering problems require solving large sparse matrix eigenproblems, where the goal is to compute only a small subset of the extreme eigenpairs, typically less than 0.5% of the matrix size. Specific applications include the discretization of partial differential equations using localized basis or grid-based methods and applications in graph theory and network analysis. The solution strategies for such eigenproblems often rely on fully iterative methods and are usually based on iterative orthogonal projection approaches. In these approaches, the large sparse matrix is orthogonally projected onto a carefully constructed smaller subspace rich in the wanted eigenvectors, followed by subspace diagonalization of the projected matrix and a subspace rotation step to recover the approximate eigenvectors of the original sparse matrix. Popular iterative approaches include Davidson [12, 13], Generalized-Davidson [14], Jacobi-Davidson [15], Chebyshev-filtered subspace iteration (ChFSI) approach [16], LOBPCG [17], and PPCG [18]. Another key class of iterative techniques is based on Krylov subspace methods, namely the Arnoldi method [19], Lanczos methods[20] and their important variants, including implicit restart Arnoldi methods [21], Krylov-Schur method [22] and block-Krylov methods [23]. Our focus in this work involves the solution of large, sparse Hermitian eigenproblems using the ChFSI approach.
@@
Chebyshev filtered subspace iteration (ChFSI) belongs to a class of iterative orthogonal projection methods and relies on a Chebyshev filtering procedure that constructs a subspace rich in the desired eigenvectors exploiting the fast growth property of Chebyshev polynomials outside the interval [-1,1], followed by a Rayleigh-Ritz step. Current work fills an important gap in the ChFSI approach by developing a residual-based approach to ChFSI that is tolerant to inexact matrix-vector products during the subspace construction step. Most importantly, we demonstrate that this approach is naturally amenable to the use of inexpensive matrix inverses while employing the ChFSI method to solve large sparse matrix generalized eigenproblems without affecting the convergence behaviour. We also show that such an approach can be robust while leveraging low-precision matrix-vector products. Such capabilities are particularly relevant given the recent modifications in today’s heterogeneous computing architectures.
@@
Modern hardware architectures have undergone substantial modifications in recent years due to the computational requirements of machine learning (ML) and artificial intelligence (AI) training. Owing to their high computational demands, these domains have necessitated the utilization of low-precision arithmetic. In response to this demand, hardware manufacturers have been enhancing support for low-precision floating-point formats, such as tensorfloat32 and bfloat16, enabling significantly faster throughput and substantial performance improvements. Recent GPU architectures exemplify this shift towards reduced-precision arithmetic. Compared to the previous Hopper architecture, NVIDIA’s Blackwell GPUs, designed for AI/ML applications, demonstrate a notable decrease in peak double-precision (FP64) floating-point performance [44]. These architectural changes underscore the necessity to modify scientific computing algorithms to efficiently utilize low-precision processes without compromising accuracy [45, 46]. A significant amount of work has been done in recent years on mixed-precision algorithms for solving linear systems of equations [47, 48, 49, 50, 51, 52, 53] using the ideas of mixed precision preconditioning and iterative refinement. Such iterative refinement ideas have also been extended to sparse eigensolvers [50, 48], but they are inefficient when we require multiple eigenpairs. For this case, the ideas of mixed precision preconditioning for the LOBPCG eigensolver [54] and mixed precision orthogonalization and Rayleigh-Ritz have been explored for the LOBPCG and Chebyshev Filtered Subspace Iteration eigensolvers [54, 38, 55]. We note that in these works the evaluation of matrix-vector products still need to be performed in the higher precision arithmetic. We further note that the capacity to accommodate approximate matrix-vector products facilitated by low-precision arithmetic can yield substantial performance enhancements. Algorithms that can accommodate approximate matrix-vector products while ensuring that errors remain within acceptable tolerances can efficiently utilize the capabilities of modern hardware, thereby reducing computational costs. The development of such algorithms becomes particularly important in light of the evolving hardware landscape.
@@
In this work, we introduce a residual-based reformulation of the Chebyshev filtered subspace iteration (ChFSI) method [56, 31, 16, 40], referred to as the R-ChFSI method, for solving large-scale sparse Hermitian eigenvalue problems. The key novelty of R-ChFSI lies in its ability to accommodate approximate matrix-vector products while preserving convergence properties. This ability makes the proposed method particularly relevant for modern hardware architectures that favor low-precision arithmetic, significantly improving computational efficiency. The proposed R-ChFSI method is also well-suited for generalized eigenvalue problems, as it can employ approximate matrix inversion and avoid expensive matrix factorizations and iterative solvers that are otherwise typically required for such problems [33]. The core of our approach is a reformulated recurrence relation that modifies the standard ChFSI update step to operate on residuals rather than the guess of the eigenvectors. We then provide a mathematical justification demonstrating that this reformulation reduces the numerical error in the Chebyshev filtered subspace construction while employing approximate matrix-vector products compared to the traditional ChFSI recurrence relation. Our analysis establishes that the proposed R-ChFSI method effectively controls error propagation in the Chebyshev filtering process, leading to faster and more reliable convergence.
@@ 2 Mathematical Background
The Chebyshev filtered subspace iteration (ChFSI) [56, 23, 31, 16, 40] approach for solving the desired eigenpairs belongs to the category of iterative orthogonal projection methods and is one of the widely used strategies to compute the smallest $n$ eigenvalues and their corresponding eigenvectors of large sparse matrices [38, 55, 57, 39, 58, 59, 60, 61, 62, 63, 64, 33]. To describe the proposed eigensolver strategy based on the ChFSI approach, in the current work, we consider the Hermitian generalized eigenvalue problem of the form
@@ 3 Approximation Methods for Chebyshev Filtered Subspace Construction
3.1 Convergence of ChFSI with approximate subspace construction
@@ 3.2 Standard Eigenvalue Problems
3.2.1 Traditional Chebyshev filtering approach employing low precision matrix-products
Employing low precision matrix-products, the recurrence relation in Eq. 3 for $\underline{\textbf{Y}}^{(i)}_{k}={C_{k}(\textbf{H})\otimes\textbf{X}^{(i)}}$, where $k=2,\dots,p$ can be written as
@@ 3.2.1 Traditional Chebyshev filtering approach employing low precision matrix-products
This is a straightforward application of the results derived in [68, section 3.5]. We note that $\gamma_{m}$ is a constant that depends on $m$, and the machine precision of the floating point arithmetic (denoted as $\varepsilon_{M}$). For the case of dense real matrices, we have $\gamma_{m}=m\varepsilon_{M}/(1-m\varepsilon_{M})$ and for dense complex matrices, it needs to be appropriately modified as described in [68, section 3.6]. Further, we also note that in the case where H is sparse, significantly tighter bounds independent of $m$ can be achieved depending on the specific implementations of the sparse matrix-dense matrix multiplication routines. ∎
@@ 3.3 Generalized Eigenvalue Problems
3.3.1 Traditional Chebyshev filtering approach with approximate matrix-products involving $\textbf{D}^{-1}$
@@
The spectral norm of the error $\bm{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}$ in the subspace construction using approximate matrix products of the recurrence relation Eq. 3 involving $\textbf{D}^{-1}$ satisfies $\lVert\bm{\Delta}_{k}^{(i)}\rVert\,\leq\zeta\eta_{k}$ for $k=0,1,\dots,p$ where $\eta_{k}$ is some finite constant that depends on $k$ and $\zeta=\lVert\textbf{D}^{-1}-\textbf{B}^{-1}\rVert\,$
@@
We now discuss this result in the context of Eq. 6. Without loss of generality, we assume that our initial guess at $i=0$ satisfies Eq. 6. While Eq. 6 is satisfied this allows for $\angle(\mathcal{S}^{(i)},\mathcal{S})$ to decrease as $i$ increases but we note that the RHS of Eq. 6 does not monotonically decrease with $\angle(\mathcal{S}^{(i)},\mathcal{S})$ and infact increases as $\angle(\mathcal{S}^{(i)},\mathcal{S})$ approaches 0. Thus, it stands to reason that beyond a certain value of $\angle(\mathcal{S}^{(i)},\mathcal{S})$ Eq. 6 no longer holds, and we cannot approach the wanted eigenspace beyond this point. Assuming that $\zeta\eta_{p}<<\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$ we estimate that the closest angle that can be achieved is
@@ 4 Results and Discussion
In the following subsections, we focus on specific problem classes, starting with the results for standard eigenvalue problems, including the cases of real symmetric and complex Hermitian matrices, followed by generalized eigenvalue problems, where we assess the convergence behaviour and the performance of the ChFSI and the R-ChFSI methods. To this end, we consider a set of three benchmark systems whose dimensions are summarized in Table 1. These eigenvalue problems are obtained from the finite-element discretization of the Kohn-Sham DFT equations for the periodic supercells of various sizes for body-centered cubic (BCC) Molybdenum systems with a single vacancy. For all the benchmarks we utilize a Chebyshev polynomial degree of 75.
@@ 4.1.1 Symmetric Eigenvalue Problems
We now consider the finite-element discretization of the Kohn-Sham DFT equations sampled at the origin (Gamma point) of the Brillouin zone for the material systems described in Table 1. This ensures that the resulting discretized equation is a real symmetric eigenvalue problem. The computation of the matrix multi-vector products with FP64 arithmetic is performed according to the methodology prescribed by [38, 55]. When using FP32 arithmetic to evaluate $\textbf{A}\textbf{R}_{\textbf{Y}}$ in Algorithm 2 of Algorithm 2, the data structures storing the matrix A and multi-vectors $\textbf{R}_{\textbf{X}}$ and $\textbf{R}_{\textbf{Y}}$ are in FP32 format and the BLAS calls used to perform $\textbf{A}\textbf{R}_{\textbf{Y}}$ are replaced with corresponding FP32 variants while the addition and scaling of multivectors is done using FP64 arithmetic. For TF32 arithmetic on NVIDIA GPUs we replace the cublasSgemm and its strided/batched variants with cublasGemmEx and its strided/batched variants with computeType set to CUBLAS_COMPUTE_32F_FAST_TF32. From Fig. 1, we observe that for all the benchmark systems summarized in Table 1, the accuracy achieved by the FP32 and TF32 variants for the R-ChFSI method is comparable to that achieved by the FP64 variant of the ChFSI method, whereas the FP32 and TF32 variants of the ChFSI method show significant reduction in achievable residual tolerance. We also note that the rate of convergence that is achieved by the FP32 and TF32 variants for the R-ChFSI method is comparable to that achieved by the FP64 variant of the ChFSI method. We also report the speedups achieved by the FP32 and TF32 variants of ChFSI and R-ChFSI over the FP64 variant of ChFSI in Fig. 2 on NVIDIA H100 GPU accelerators. We note that we obtain similar speedups for both the ChFSI and R-ChFSI methods allowing us to conclude that the additional overheads in the R-ChFSI method are negligible. We have achieved speedups of up to 1.75x for the TF32 variant of the R-ChFSI method over the baseline implementation of the ChFSI method with FP64 arithmetic. We also report the speedups achieved by the FP32 variant of ChFSI and R-ChFSI over the FP64 variant of ChFSI in Fig. 3 on Intel Xeon Platinum 8480C CPUs. We note that the speedups of the R-ChFSI method are slightly lower than ChFSI method due to the additional overheads in the R-ChFSI method. We achieve speedups of up to 2.1x for the FP32 variant of the R-ChFSI method over the baseline implementation of the ChFSI method with FP64 arithmetic.
@@ 4.1.2 Hermitian Eigenvalue Problems
We now consider the finite-element discretization of the Kohn-Sham DFT equations at a non-zero $k$-point  [38] in the Brillouin zone for the systems described in Table 1. The resulting discretized equation is now a complex Hermitian eigenvalue problem. The computation of the matrix multi-vector products with FP64 arithmetic is done according to the methodology prescribed by [38, 55]. When using FP32 arithmetic to evaluate $\textbf{A}\textbf{R}_{\textbf{Y}}$ in Algorithm 2 of Algorithm 2, the data structures storing the matrix A and multi-vectors $\textbf{R}_{\textbf{X}}$ and $\textbf{R}_{\textbf{Y}}$ are in FP32 format and the BLAS calls used to perform $\textbf{A}\textbf{R}_{\textbf{Y}}$ are replaced with corresponding FP32 variants while the additon and scaling of multivectors is done using FP64 arithmetic. For TF32 arithmetic on NVIDIA GPUs we replace the cublasCgemm and its strided/batched variants with cublasGemmEx and its strided/batched variants with computeType set to CUBLAS_COMPUTE_32F_FAST_TF32.

### v2

@@ 1 Introduction
Many scientific and engineering problems require solving large sparse matrix eigenproblems, where the goal is to compute only a small subset of the extreme eigenpairs, typically less than 0.5% of the matrix size. Specific applications include the discretization of partial differential equations using localized basis or grid-based methods and applications in graph theory and network analysis. The solution strategies for such eigenproblems often rely on fully iterative methods and are usually based on iterative orthogonal projection approaches. In these approaches, the large sparse matrix is orthogonally projected onto a carefully constructed smaller subspace rich in the wanted eigenvectors, followed by subspace diagonalization of the projected matrix and a subspace rotation step to recover the eigenvector estimates of the original sparse matrix. Popular iterative approaches include Davidson [12, 13], Generalized-Davidson [14], Jacobi-Davidson [15], Chebyshev-filtered subspace iteration (ChFSI) approach [16], LOBPCG [17], and PPCG [18]. Another key class of iterative techniques is based on Krylov subspace methods, namely the Arnoldi method [19], Lanczos methods[20] and their important variants, including implicit restart Arnoldi methods [21], Krylov-Schur method [22] and block-Krylov methods [23]. Our focus in this work involves the solution of large, sparse Hermitian eigenproblems using the ChFSI approach.
@@
Chebyshev filtered subspace iteration (ChFSI) belongs to a class of iterative orthogonal projection methods and relies on a Chebyshev filtering procedure that constructs a subspace rich in the desired eigenvectors exploiting the fast growth property of Chebyshev polynomials outside the interval [-1,1], followed by a Rayleigh-Ritz step. Current work fills an important gap in the ChFSI approach by developing a residual-based approach to ChFSI that is tolerant to inexact matrix-vector products during the subspace construction step. Most importantly, we demonstrate that this approach is naturally amenable to the use of inexpensive matrix inverses while employing the ChFSI method to solve large sparse matrix generalized eigenproblems without affecting the convergence behavior. We also show that such an approach can be robust while leveraging low-precision matrix-vector products. Such capabilities are particularly relevant given the recent modifications in today’s heterogeneous computing architectures.
@@
Modern hardware architectures have undergone substantial modifications in recent years due to the computational requirements of machine learning (ML) and artificial intelligence (AI) training. Owing to their high computational demands, these domains have necessitated the utilization of low-precision arithmetic. In response to this demand, hardware manufacturers have been enhancing support for low-precision floating-point formats, such as tensorfloat32 and bfloat16, enabling significantly faster throughput and substantial performance improvements. Recent GPU architectures exemplify this shift towards reduced-precision arithmetic. Compared to the previous Hopper architecture, NVIDIA’s Blackwell GPUs, designed for AI/ML applications, demonstrate a notable decrease in peak double-precision (FP64) floating-point performance [44]. These architectural changes underscore the necessity to modify scientific computing algorithms to efficiently utilize low-precision processes without compromising accuracy [45, 46]. A significant amount of work has been done in recent years on mixed-precision algorithms for solving linear systems of equations [47, 48, 49, 50, 51, 52, 53] using the ideas of mixed precision preconditioning and iterative refinement. Such iterative refinement ideas have also been extended to sparse eigensolvers [50, 48], but they are inefficient when we require multiple eigenpairs. For this case, the ideas of mixed precision preconditioning for the LOBPCG eigensolver [54] and mixed precision orthogonalization and Rayleigh-Ritz have been explored for the LOBPCG and Chebyshev Filtered Subspace Iteration eigensolvers [54, 38, 55]. We note that in these works the evaluation of matrix-vector products still need to be performed in the higher precision arithmetic. We further note that the capacity to accommodate inexact matrix-vector products facilitated by low-precision arithmetic can yield substantial performance enhancements. Algorithms that can accommodate inexact matrix-vector products while ensuring that errors remain within acceptable tolerances can efficiently utilize the capabilities of modern hardware, thereby reducing computational costs. The development of such algorithms becomes particularly important in light of the evolving hardware landscape.
@@
In this work, we introduce a residual-based reformulation of the Chebyshev filtered subspace iteration (ChFSI) method [56, 31, 16, 40], referred to as the R-ChFSI method, for solving large-scale sparse Hermitian eigenvalue problems. The key novelty of R-ChFSI lies in its ability to accommodate inexact matrix-vector products while preserving convergence properties. This ability makes the proposed method particularly relevant for modern hardware architectures that favor low-precision arithmetic, significantly improving computational efficiency. The proposed R-ChFSI method is also well-suited for generalized eigenvalue problems, as it can employ approximate matrix inversion and avoid expensive matrix factorizations and iterative solvers that are otherwise typically required for such problems [33]. The core of our approach is a reformulated recurrence relation that modifies the standard ChFSI update step to operate on residuals rather than the guess of the eigenvectors. We then provide a mathematical justification demonstrating that this reformulation reduces the numerical error in the Chebyshev filtered subspace construction while employing inexact matrix-vector products compared to the traditional ChFSI recurrence relation. Our analysis establishes that the proposed R-ChFSI method effectively controls error propagation in the Chebyshev filtering process, leading to faster and more reliable convergence.
@@ 2 Mathematical Background
The Chebyshev filtered subspace iteration (ChFSI) [56, 23, 31, 16, 40] approach for solving the desired eigenpairs belongs to the category of iterative orthogonal projection methods and is one of the widely used strategies to compute the smallest $n$ eigenvalues and their corresponding eigenvectors of large sparse matrices [57, 38, 55, 58, 39, 59, 60, 61, 62, 63, 64, 65, 33]. To describe the proposed eigensolver strategy based on the ChFSI approach, in the current work, we consider the Hermitian generalized eigenvalue problem of the form
@@ 3 Approximation Methods for Chebyshev Filtered Subspace Construction
3.1 Convergence of ChFSI with inexact subspace construction
@@ 3.2 Standard Eigenvalue Problems
3.2.1 Traditional Chebyshev filtering approach employing low-precision matrix-products
Employing low-precision matrix-products, the recurrence relation in Eq. 3 for $\underline{\textbf{Y}}^{(i)}_{k}={C_{k}(\textbf{H})\otimes\textbf{X}^{(i)}}$, where $k=2,\dots,p$ can be written as
@@ 3.2.1 Traditional Chebyshev filtering approach employing low-precision matrix-products
This is a straightforward application of the results derived in [69, section 3.5]. We note that $\gamma_{m}$ is a constant that depends on $m$, and the machine precision of the floating point arithmetic (denoted as $\varepsilon_{M}$). For the case of dense real matrices, we have $\gamma_{m}=m\varepsilon_{M}/(1-m\varepsilon_{M})$ and for dense complex matrices, it needs to be appropriately modified as described in [69, section 3.6]. Further, we also note that in the case where H is sparse, significantly tighter bounds independent of $m$ can be achieved depending on the specific implementations of the sparse matrix-dense matrix multiplication routines. ∎
@@ 3.3 Generalized Eigenvalue Problems
3.3.1 Traditional Chebyshev filtering approach with inexact matrix-products involving $\textbf{D}^{-1}$
@@
The spectral norm of the error $\bm{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}$ in the subspace construction using inexact matrix products of the recurrence relation Eq. 3 involving $\textbf{D}^{-1}$ satisfies $\lVert\bm{\Delta}_{k}^{(i)}\rVert\,\leq\zeta\eta_{k}$ for $k=0,1,\dots,p$ where $\eta_{k}$ is some finite constant that depends on $k$ and $\zeta=\lVert\textbf{D}^{-1}-\textbf{B}^{-1}\rVert\,$
@@
We now discuss this result in the context of Eq. 6. Without loss of generality, we assume that our initial guess at $i=0$ satisfies Eq. 6. While Eq. 6 is satisfied this allows for $\angle(\mathcal{S}^{(i)},\mathcal{S})$ to decrease as $i$ increases, but we note that the RHS of Eq. 6 does not monotonically decrease with $\angle(\mathcal{S}^{(i)},\mathcal{S})$ and in fact increases as $\angle(\mathcal{S}^{(i)},\mathcal{S})$ approaches 0. Thus, it stands to reason that beyond a certain value of $\angle(\mathcal{S}^{(i)},\mathcal{S})$ Eq. 6 no longer holds, and we cannot approach the wanted eigenspace beyond this point. Assuming that $\zeta\eta_{p}<<\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$ we estimate that the closest angle that can be achieved is
@@ 4 Results and Discussion
In the following subsections, we focus on specific problem classes, starting with the results for standard eigenvalue problems, including the cases of real symmetric and complex Hermitian matrices, followed by generalized eigenvalue problems, where we assess the convergence behavior and the performance of the ChFSI and the R-ChFSI methods. To this end, we consider a set of three benchmark systems whose dimensions are summarized in Table 1. These eigenvalue problems are obtained from the finite-element discretization of the Kohn-Sham DFT equations for the periodic supercells of various sizes for body-centered cubic (BCC) Molybdenum systems with a single vacancy. For all the benchmarks we utilize a Chebyshev polynomial degree of 75.
@@ 4.1.1 Symmetric Eigenvalue Problems
We now consider the finite-element discretization of the Kohn-Sham DFT equations sampled at the origin (Gamma point) of the Brillouin zone for the material systems described in Table 1. This ensures that the resulting discretized equation is a real symmetric eigenvalue problem. The computation of the matrix multi-vector products with FP64 arithmetic is performed according to the methodology prescribed by [38, 55]. When using FP32 arithmetic to evaluate $\textbf{A}\textbf{R}_{\textbf{Y}}$ in Algorithm 2 of Algorithm 2, the data structures storing the matrix A and multi-vectors $\textbf{R}_{\textbf{X}}$ and $\textbf{R}_{\textbf{Y}}$ are in FP32 format and the BLAS calls used to perform $\textbf{A}\textbf{R}_{\textbf{Y}}$ are replaced with corresponding FP32 variants while the addition and scaling of multi-vectors is done using FP64 arithmetic. For TF32 arithmetic on NVIDIA GPUs we replace the cublasSgemm and its strided/batched variants with cublasGemmEx and its strided/batched variants with computeType set to CUBLAS_COMPUTE_32F_FAST_TF32. From Fig. 1, we observe that for all the benchmark systems summarized in Table 1, the accuracy achieved by the FP32 and TF32 variants for the R-ChFSI method is comparable to that achieved by the FP64 variant of the ChFSI method, whereas the FP32 and TF32 variants of the ChFSI method show significant reduction in achievable residual tolerance. We also note that the rate of convergence that is achieved by the FP32 and TF32 variants for the R-ChFSI method is comparable to that achieved by the FP64 variant of the ChFSI method. We also report the speedups achieved by the FP32 and TF32 variants of ChFSI and R-ChFSI over the FP64 variant of ChFSI in Fig. 2 on NVIDIA H100 GPU accelerators. We note that we obtain similar speedups for both the ChFSI and R-ChFSI methods allowing us to conclude that the additional overheads in the R-ChFSI method are negligible. We have achieved speedups of up to 1.75x for the TF32 variant of the R-ChFSI method over the baseline implementation of the ChFSI method with FP64 arithmetic. We also report the speedups achieved by the FP32 variant of ChFSI and R-ChFSI over the FP64 variant of ChFSI in Fig. 3 on Intel Xeon Platinum 8480C CPUs. We note that the speedups of the R-ChFSI method are slightly lower than ChFSI method due to the additional overheads in the R-ChFSI method. We achieve speedups of up to 2.1x for the FP32 variant of the R-ChFSI method over the baseline implementation of the ChFSI method with FP64 arithmetic.
@@ 4.1.2 Hermitian Eigenvalue Problems
We now consider the finite-element discretization of the Kohn-Sham DFT equations at a non-zero $k$-point  [38] in the Brillouin zone for the systems described in Table 1. The resulting discretized equation is now a complex Hermitian eigenvalue problem. The computation of the matrix multi-vector products with FP64 arithmetic is done according to the methodology prescribed by [38, 55]. When using FP32 arithmetic to evaluate $\textbf{A}\textbf{R}_{\textbf{Y}}$ in Algorithm 2 of Algorithm 2, the data structures storing the matrix A and multi-vectors $\textbf{R}_{\textbf{X}}$ and $\textbf{R}_{\textbf{Y}}$ are in FP32 format and the BLAS calls used to perform $\textbf{A}\textbf{R}_{\textbf{Y}}$ are replaced with corresponding FP32 variants while the addition and scaling of multi-vectors is done using FP64 arithmetic. For TF32 arithmetic on NVIDIA GPUs we replace the cublasCgemm and its strided/batched variants with cublasGemmEx and its strided/batched variants with computeType set to CUBLAS_COMPUTE_32F_FAST_TF32.